OpenStudy (anonymous):

Raise the quantity in parenthesis to the indicated exponent, and simplify the resulting expression.Express answers with positive exponents. (-36x^2y/180x^-3y^-5)^-2 This is as far as I got if it is even correct so far.... (y^6/3x^2)^-2 ????? but if this is right I'm not sure of what to do from here.

OpenStudy (lgbasallote):

uhh i dont think that's rgiht...

OpenStudy (lgbasallote):

\[-\frac{36}{180} = -\frac 15\] \[\large \frac{x^2}{x^{-3}} = x^5\] \[\frac{y}{y^{-5}} = y^6\]

OpenStudy (anonymous):

you know what? i had wrote on my paper that 180 divided by 36 was five, but I guess I overlooked that when I was doing my problem. thanks:)

OpenStudy (lgbasallote):

haha welcome ^)^

OpenStudy (anonymous):

I have a question about x^2/x^-3 = 5 ... are we adding them? if so how does it become positive?

OpenStudy (anonymous):

and where did the ^-2 go that was in parenthesis around the problem?

OpenStudy (lgbasallote):

\[\Large \frac{x^2}{x^{-3}} \implies x^{2 - (-3)} \implies x^{2 + 3} \implies x^5\]

OpenStudy (lgbasallote):

and the ^-2 is still there...i was just pointing out your mistakes

OpenStudy (anonymous):

oh ok thanks